# Gradient Attention Map Based Verification of Deep Convolutional Neural Networks with Application to X-ray Image Datasets

### v1

@@ 1 Introduction
In orthodontics, precise classification is crucial for determining the appropriate treatment pathway. For example, in Class III malocclusion diagnosis, lateral cephalometric X-rays and profile photographs are commonly used to classify whether a patient requires surgical intervention [6, 7]. Similarly, skeletal maturation assessment using cranial base synchondroses, such as the spheno-occipital synchondrosis (SOS), is vital for forensic age estimation and orthodontic treatment planning [1, 2]. In both applications, misclassification can lead to incorrect diagnoses and suboptimal treatment decisions, emphasizing the need for reliable models.
Despite advancements in deep learning, model misapplication remains a critical issue. If a model trained on malocclusion classification is mistakenly tested on SOS data, it would likely generate incorrect predictions due to the fundamental differences in anatomical structures and classification objectives. Traditional evaluation methods, such as accuracy, precision, and recall, do not provide insight into whether the model was trained on a compatible dataset.

### v2

@@ 1 Introduction
In orthodontics, precise classification is crucial for determining the appropriate treatment pathway. For example, in Class III malocclusion diagnosis, lateral cephalometric X-rays and profile photographs are commonly used to classify whether a patient requires surgical intervention [6, 7]. Similarly, skeletal maturation assessment using cranial base synchondroses, such as the spheno-occipital synchondrosis (SOS), is vital for forensic age estimation and orthodontic treatment planning [1, 2]. In both applications, misclassification can lead to incorrect diagnoses and suboptimal treatment decisions, emphasizing the need for reliable models. Despite advancements in deep learning, model misapplication remains a critical issue. For example, if a model trained on malocclusion classification is mistakenly tested on SOS data, it may generate incorrect predictions due to fundamental differences in anatomical structures and classification objectives. Traditional evaluation metrics, such as accuracy, precision, and recall, do not provide insight into whether the model was trained on a compatible dataset.
To ensure consistency and clinical relevance, our model selection and experimental setup build on previously validated deep learning frameworks for orthodontic diagnosis and treatment planning. Specifically, we utilize models developed in prior work on third molar development classification [22] and treatment decision-making for anterior open bite cases [23], both of which demonstrate robust performance on cephalometric, CBCT, and tabular data.